# Discrete flavour and CP symmetries in light of JUNO and neutrino global fit

### v1

@@ 4.1 The breaking patterns of $A_{5}\rtimes H_{CP}$
We see that the mixing parameters depend on the continuous parameters $\theta_{\ell}$ and $\theta_{\nu}$. As a consequence, sum rules among the mixing angles and the Dirac CP phase $\delta_{CP}$ can be found as follows:
$U^{VI}_{1},U^{VI}_{2}$~ $:$ $~~~\cos^{2}\theta_{12}\cos^{2}\theta_{13}=\frac{\phi_{g}^{2}}{4}\,$,
@@ 5 Numerical analysis
In the following, we will first confront the predictions of the one- and two-parameter mixing patterns discussed in sections 3 and 4 with the latest NuFIT results. A given lepton mixing pattern is regarded as phenomenologically viable if its predictions for the mixing parameters at the minimum of the $\chi^{2}$ function lie within the corresponding $3\sigma$ intervals summarized in table 3. After that, we confront the predicted values of $\sin^{2}\theta_{12}$ with the most recent 59.1-day JUNO measurement, as well as with the projected precision expected after 6 years of JUNO data taking. We also examine the predictions for $\sin^{2}\theta_{23}$ and $\delta_{CP}$ in light of the anticipated sensitivity reaches of the future long-baseline experiments DUNE and T2HK. Among the one- and two-parameter mixing patterns, several patterns—namely $U^{I}$, $U^{II}$, $U^{VIII}$, $U^{XI}$ and $U^{X}$—arise from the series of flavour groups $\Delta(6n^{2})\rtimes H_{CP}$ (with $U^{VIII}$ also obtainable from $D_{n}\rtimes H_{CP}$). To keep the analysis tractable and avoid excessively large group orders, we restrict our attention to groups $\Delta(6n^{2})$ with $n\leq 4$ and $D_{n}$ with $n\leq 8$ in what follows.
@@ 5.1 Lepton mixing matrix depending on one free parameter
For the one-parameter mixing patterns, $U^{I}$, $U^{II}$, $U^{III}$, $U^{IV}$ and $U^{V}$, the best-fit value of the free parameter $\theta$, the minimum $\chi^{2}_{\text{min}}$ and the resulting predictions for the lepton mixing angles $\theta_{12}$, $\theta_{13}$, $\theta_{23}$ as well as the CPV phases $\delta_{CP}$, $\alpha_{21}$ and $\alpha_{31}$ for both NO and IO neutrino mass spectra are collected in table 4. Note that the entries in the $\sin^{2}\theta_{12}$ column highlighted in red indicate predictions that fall within the $3\sigma$ interval of the JUNO 59.1-day measurement [99],
@@
Confronting the results in table 4 with the NuFIT $3\sigma$ allowed regions given in table 3, we find that a total of 8 (6) one-parameter mixing patterns are phenomenologically viable for NO (IO), once all inequivalent row and column permutations of the lepton mixing matrix are taken into account. For the mixing patterns $U^{I}$ and $U^{II}$ arising from $\Delta(6n^{2})\rtimes H_{CP}$, the smallest group that can reproduce the NuFIT data is $\Delta(24)\cong S_{4}$ with $n=2$. In case of $(\varphi_{1},\varphi_{2})=\left(-\frac{\pi}{2},\,\frac{\pi}{2}\right)$, the pattern $U^{I}$ reduces to the TM1 form [100, 101] with its first column fixed as $\left(-\frac{2}{\sqrt{6}},\,-\frac{1}{\sqrt{6}},\,\frac{1}{\sqrt{6}}\right)^{T}$, as can seen from Eq. (3.3), and it agrees with the NuFIT limits for both NO and IO. The TM2-type patterns, namely $U^{II}$ with $(\varphi_{1},\varphi_{2})=(0,\pi)$ and $U^{V}$ featuring the trimaximal column $\left(\frac{1}{\sqrt{3}},\,\frac{1}{\sqrt{3}},\,\frac{1}{\sqrt{3}}\right)^{T}$ (see Eq. (3.15) and Eq. (3.37)), don’t agree with JUNO’s latest results at $3\sigma$ level, although it is very close to the $3\sigma$ upper limit of NuFIT for both NO and IO. Among the one-parameter patterns originating from $A_{5}\rtimes H_{CP}$, the GR2 structure $U^{III}$—which preserves the second column of the Golden Ratio mixing pattern, $\left(\sqrt{\frac{1}{\sqrt{5}\,\phi_{g}}},\,\sqrt{\frac{\phi_{g}}{2\sqrt{5}}},\,\sqrt{\frac{\phi_{g}}{2\sqrt{5}}}\right)^{T}$ (see Eq. (3.26))—remains compatible with the NuFIT constraints in both NO and IO. By contrast, the real pattern $U^{IV}$, characterized by the fixed first column $\left(\frac{\phi_{g}}{2},\,\frac{\phi_{g}-1}{2},\,\frac{1}{2}\right)^{T}$ (see Eq. (3.32)), is only viable in NO.
@@
We display in figure 1 the best-fit values of the three mixing angles and $\delta_{CP}$ for all viable one-parameter models under both NO and IO neutrino mass spectrum. The dashed lines indicate the central values of the observables from the NuFIT global fit. The light-red band in the $\sin^{2}\theta_{12}$ panel corresponds to the $3\sigma$ interval derived from the recent JUNO measurement, see Eq. (5.4), while the darker and narrower red band indicates the projected $3\sigma$ precision after six years of data taking, assuming the current JUNO central value $\sin^{2}\theta_{12}=0.3092$ and an expected uncertainty of about $0.5\%$ [31]. For NO, three one-parameter mixing patterns — $U^{I}$ (TM1), $U^{IV}_{1}$ and $U^{IV}_{2}$ — yield predictions for $\sin^{2}\theta_{12}$ that fall within the current JUNO constraint. In contrast, the patterns $U^{II}$ (TM2), $U^{III}$ (GR2), and $U^{V}$ (TM2) are disfavored by the JUNO data at the $3\sigma$ level. None of the one-parameter patterns, however, would remain compatible with the much better sensitivity expected after six years of operation if the current best fit value would not change significantly. For IO, only the pattern $U^{I}$ (TM1) lies in the present JUNO $3\sigma$ allowed interval, and all patterns would be excluded once the planned future precision is reached. We also observe that the predicted values of $\sin^{2}\theta_{13}$ for nearly all patterns lie very close to the NuFIT best-fit line, reflecting the excellent experimental precision already achieved for this observable. Finally, the red bands shown for $\sin^{2}\theta_{23}$ and $\delta_{CP}$ represent the anticipated future $3\sigma$ sensitivities of next-generation long-baseline experiments such as DUNE [92] and T2HK [93]. It is worth noting that the future experimental uncertainties on $\sin^{2}\theta_{23}$ and $\delta_{CP}$ depend sensitively on their true values. For $\sin^{2}\theta_{23}$, as illustrated in Ref. [102], the projected precision of DUNE and T2HK ranges from about $1\%$ when the true value lies near the edges of the current $3\sigma$ interval, to roughly $7\%$ around the maximal-mixing point $\sin^{2}\theta_{23}=0.5$. For the present NuFIT best-fit values, $\sin^{2}\theta_{23}=0.47$ (NO) and $\sin^{2}\theta_{23}=0.55$ (IO), the expected uncertainty is approximately $3\%$, which we adopt as a representative estimate. For $\delta_{CP}$, the projected $1\sigma$ precision of DUNE is approximately $6^{\circ}$ if the true value is $\delta_{CP}=0^{\circ}$, while it degrades to about $16^{\circ}$ for $\delta_{CP}=270^{\circ}$. In case of T2HK, the corresponding $1\sigma$ uncertainties are around $6^{\circ}$ for $\delta_{CP}=0^{\circ}$ and roughly $20^{\circ}$ for $\delta_{CP}=270^{\circ}$. Taking these variations into account, we adopt $12^{\circ}$ as the expected $1\sigma$ precision for $\delta_{CP}$ in our analysis for illustration. Combining the numerical results in table 4, we observe that the patterns $U^{I}$, $U^{II}$, and $U^{III}$ predict maximal atmospheric mixing $\theta_{23}$ and Dirac CPV phase $\delta_{CP}$. In contrast, the real patterns $U^{IV}_{1}$ and $U^{IV}_{2}$ imply CP conservation. Neither maximal nor vanishing $\delta_{CP}$ is achievable within $U^{V}_{1}$ and $U^{V}_{2}$. Hence precise measurements of $\theta_{23}$ and $\delta_{CP}$ will therefore play a important role in discriminating among these viable mixing schemes.
@@
For $\sin^{2}\theta_{12}$, the predicted profiles are exceedingly narrow because $\sin^{2}\theta_{12}$ is tightly constrained by the small uncertainties in $\sin^{2}\theta_{13}$, as implied by Eqs. (3.8), (3.29), (3.35) and (3.42) for $U^{I}$, $U^{III}$, $U^{IV}_{1}$ and $U^{V}_{1}$, respectively. In the NO case, the prediction of $U^{V}_{1}$ for $\sin^{2}\theta_{12}$ is disfavored by the latest JUNO data, whereas $U^{I}$ and $U^{IV}_{1}$ remain viable at $3\sigma$. In the IO case, $U^{III}$ is excluded by current JUNO data, while $U^{I}$ stays consistent. If JUNO achieves the projected $0.5\%$ precision and the current best fit values would not change significantly, all these patterns could be tested in both NO and IO cases. These conclusions regarding $\sin^{2}\theta_{12}$ for the one-parameter mixing patterns are fully consistent with the earlier analysis based on the best-fit results. For $\sin^{2}\theta_{23}$, the patterns $U^{I}$ and $U^{III}$ predict the fixed value $\sin^{2}\theta_{23}=0.5$, while $U^{IV}_{1}$ and $U^{V}_{1}$ yield sharply peaked predictions with $\sin^{2}\theta_{23}>0.5$, due to correlations with $\sin^{2}\theta_{13}$ encoded in Eqs. (3.35) and (3.42). Taking future sensitivities into account, the predicted ranges of $U^{IV}_{1}$ and $U^{V}_{1}$ could be probed in the NO case, while in the IO case both $U^{I}$ and $U^{III}$ could be tested. Regarding the Dirac phase, $U^{I}$ and $U^{III}$ predict $\delta_{CP}=1.5\pi$, while $U^{IV}_{1}$ predicts $\delta_{CP}=\pi$. i.e., absence of CP-violation in neutrino oscillations. In the NO case, the prediction $\delta_{CP}=1.5\pi$ lies outside the projected $3\sigma$ allowed range of the prospective DUNE and T2HK data, whereas $\delta_{CP}=\pi$ remains within it. In contrast, for the IO case, $\delta_{CP}=1.5\pi$ falls within the $3\sigma$ allowed interval.
@@
In summary, the JUNO 59.1-day data imposes significant restrictions on the landscape of one-parameter mixing patterns. The $3\sigma$ interval of $\sin^{2}\theta_{12}$ from JUNO excludes several patterns that would otherwise remain compatible with the NuFIT global fit. We have focused on the flavour symmetry groups $A_{5}$, $\Sigma(168)$ and $D(6n^{2})$ with $n\leq 4$. For NO, we find that only $U^{I}$ (TM1), originating from $S_{4}\rtimes H_{CP}$, and $U^{IV}_{1,2}$, originating from $A_{5}\rtimes H_{CP}$, predict $\sin^{2}\theta_{12}$ values consistent with the current JUNO constraint. For IO, the viable space shrinks further, leaving $U^{I}$ (TM1) as the sole surviving option. Patterns such as $U^{II}$ (TM2), $U^{III}$ (GR2) and $U^{V}$ (TM2), although acceptable under NuFIT alone, become disfavored once the JUNO likelihood is taken into account. It should be emphasized that, for the mixing pattern $U^{I}$, additional phenomenologically viable textures—compatible with the current JUNO data—will emerge other than the TM1 form once higher values of $n$ in the group $\Delta(6n^{2})$ are considered. The future precision of $\sin^{2}\theta_{12}$, $\sin^{2}\theta_{23}$, and $\delta_{CP}$ will provide additional discriminatory power among these one-parameter mixing patterns.
@@ 5.2 Lepton mixing matrix depending on two free parameters
For the two-parameter mixing patterns derived from $A_{5}\rtimes H_{CP}$, we find 10 and 4 phenomenologically viable realizations of $U^{VI}$ for NO and IO, respectively, as listed in table 5. The patterns $U^{VI}_{2}$, $U^{VI}_{4,1}$, $U^{VI}_{4,2}$, $U^{VI}_{4,3}$, $U^{VI}_{4,4}$, $U^{VI}_{6,1}$ and $U^{VI}_{6,2}$ predict CP-conserving Dirac and Majorana phases, consistent with the general argument presented in section 4.1. Consequently, these patterns are compatible with present data in the NO case, while they are disfavored for IO mass spectrum. By contrast, $U^{VI}_{1}$, $U^{VI}_{3,2}$ and $U^{VI}_{5}$ yield nontrivial CPV phases and can accommodate the experimental data for both NO and IO. $U^{VI}_{3,1}$ can provide viable predictions only for the IO scenario. Under the sum rule constraints in Eq. (4.7), the allowed regions of the lepton observables given in table 10 become narrower than the corresponding $3\sigma$ intervals in table 3. The likelihood profiles for $\sin^{2}\theta_{12}$, $\sin^{2}\theta_{23}$ and $\delta_{CP}$ are shown in figure 4 for selected viable patterns $U^{VI}$ with $\chi^{2}_{\min}\leq 9$. The patterns $U^{VI}_{4,1}$ and $U^{VI}_{6,2}$ for NO, as well as $U^{VI}_{3,1}$, $U^{VI}_{3,2}$ and $U^{VI}_{5}$ for IO, are not shown in the figure.
@@
For $\sin^{2}\theta_{12}$, figure 4 shows that the patterns $U^{VI}_{4,2}$, $U^{VI}_{4,4}$, and $U^{VI}_{6,1}$ are consistent with the NuFIT data at the $2\sigma$ level, whereas $U^{VI}_{1}$, $U^{VI}_{2}$, $U^{VI}_{3,2}$, $U^{VI}_{4,3}$, and $U^{VI}_{5}$ remain compatible at $3\sigma$ for the NO scenario. In case of IO, only $U^{VI}_{1}$ is consistent at $2\sigma$. The patterns $U^{VI}_{1}$ and $U^{VI}_{2}$ yield very sharp predictions for $\sin^{2}\theta_{12}$, since $\sin^{2}\theta_{12}$ is correlated with $\sin^{2}\theta_{13}$, see Eq. (4.7). The predictions of the patterns $U^{VI}_{1}$, $U^{VI}_{2}$, $U^{VI}_{3,2}$, and $U^{VI}_{5}$ partially overlap with the $3\sigma$ range of the JUNO 59.1-day measurement, but they are disfavored at the $2\sigma$ level and are expected to be excluded once JUNO reaches its projected six-year precision of $0.5\%$. The remaining four cases $U^{VI}_{4,2}$, $U^{VI}_{4,3}$, $U^{VI}_{4,4}$ and $U^{VI}_{6,1}$ are consistent with both the current and projected JUNO sensitivities for NO and IO.
@@
For $\sin^{2}\theta_{23}$, the patterns $U^{VI}_{3,2}$, $U^{VI}_{4,2}$, $U^{VI}_{4,3}$ and $U^{VI}_{5}$ could be excluded by the projected DUNE and T2HK sensitivities in the NO case, as their predicted values of $\theta_{23}$ lie in second octant. For IO, the pattern $U^{VI}_{1}$ remains consistent with the anticipated experimental precision. Moreover, all patterns that predict CP-conserving Dirac phase can be tested by future measurements, whereas those yielding nontrivial $\delta_{CP}$ remain viable under the projected sensitivities.
@@
For the two-parameter mixing patterns originating from $\Sigma(168)\rtimes H_{CP}$, we identify 5 viable realizations for NO and 2 for IO, as summarized in table 5. In the NO case, the four patterns $U^{VII}_{2,i}$ with $i=1,2,3,4$ predict trivial Dirac phase, since $U^{VII}_{2,i}$ is real matrix as follows from Eq. (4.14). The pattern $U^{VII}_{1,1}$ instead yields CP-violating Dirac phase and it remains compatible with global oscillation data for both mass orderings, whereas $U^{VII}_{1,2}$ is viable only for IO. Figure 4 displays the likelihood profiles of 3 cases $U^{VII}_{2,2}$, $U^{VII}_{2,3}$ and $U^{VII}_{2,4}$. All of them predict $\sin^{2}\theta_{12}$ values lying within both the current JUNO 59.1-day and the projected six-year precision $3\sigma$ intervals. Their predictions for $\sin^{2}\theta_{23}$, however, show different levels of compatibility with future accelerator experiments sensitivities provided the current NuFIT best fit value including the SK data would not change drastically: while the $U^{VII}_{2,2}$ and $U^{VII}_{2,4}$ predictions fall outside the expected DUNE and T2HK ranges, the allowed region of $U^{VII}_{2,3}$ partially overlaps with the projected sensitivities. Since all these patterns exhibit CP-conserving Dirac phase, they will be critically tested once DUNE and T2HK achieve their anticipated precision on $\delta_{CP}$.

### v2

@@ 4.1 The breaking patterns of $A_{5}\rtimes H_{CP}$
We see that the mixing parameters depend on the continuous parameters $\theta_{\ell}$ and $\theta_{\nu}$. As a consequence, sum rules among the mixing angles and the Dirac CP phase $\delta_{CP}$ can be found as follows  [^9] The sum rule relating $\cos^{2}\theta_{12}$ and $\cos^{2}\theta_{13}$ in cases $U^{VI}_{1}$ and $U^{VI}_{2}$ coincides with the sum rule in Eq. (3.34) since $\phi_{g}^{2}=1+\phi_{g}$. It holds also in an $A_{5}$ model with no gCP symmetry having three real parameters [56], in which $\sin^{2}\theta_{23}$ and $\delta_{CP}$ are unconstrained [50].:
$U^{VI}_{1},U^{VI}_{2}$~ $:$ $~~~4\cos^{2}\theta_{12}\cos^{2}\theta_{13}=\phi_{g}^{2}\,$,
@@ 5 Numerical analysis
In the following, we will first confront the predictions of the one- and two-parameter mixing patterns discussed in sections 3 and 4 with the latest NuFIT results. A given lepton mixing pattern is regarded as phenomenologically viable if its predictions for the mixing parameters at the minimum of the $\chi^{2}$ function lie within the corresponding $3\sigma$ intervals summarized in table 3. After that, we confront the predicted values of $\sin^{2}\theta_{12}$ with the most recent 59.1-day JUNO measurement, as well as with the projected precision expected after 6 years of JUNO data taking. We also examine the predictions for $\sin^{2}\theta_{23}$ and $\delta_{CP}$ in light of the anticipated sensitivity reaches of the future long-baseline experiments DUNE and T2HK.
[…]
Finally, we note that among the one- and two-parameter mixing patterns, several patterns—namely $U^{I}$, $U^{II}$, $U^{VIII}$, $U^{XI}$ and $U^{X}$—arise from the series of flavour groups $\Delta(6n^{2})\rtimes H_{CP}$ (with $U^{VIII}$ also obtainable from $D_{n}\rtimes H_{CP}$). To keep the analysis tractable and avoid excessively large group orders, we restrict our attention to groups $\Delta(6n^{2})$ with $n\leq 4$ and $D_{n}$ with $n\leq 8$ in what follows.
@@ 5.1 Lepton mixing matrix depending on one free parameter
For the one-parameter mixing patterns, $U^{I}$, $U^{II}$, $U^{III}$, $U^{IV}$ and $U^{V}$, the best-fit value of the free parameter $\theta$, the minimum $\chi^{2}_{\text{min}}$ and the resulting predictions for the lepton mixing angles $\theta_{12}$, $\theta_{13}$, $\theta_{23}$ as well as the CPV phases $\delta_{CP}$, $\alpha_{21}$ and $\alpha_{31}$ for both NO and IO neutrino mass spectra are collected in table 4. We note that the entries in the $\sin^{2}\theta_{12}$ column highlighted in red indicate predictions that fall within the $3\sigma$ interval of the JUNO 59.1-day measurement [33],
@@
Confronting the results in table 4 with the NuFIT $3\sigma$ allowed regions given in table 3, we find that a total of 8 (6) one-parameter mixing patterns are phenomenologically viable for NO (IO), once all inequivalent row and column permutations of the lepton mixing matrix are taken into account. For the mixing patterns $U^{I}$ and $U^{II}$ arising from $\Delta(6n^{2})\rtimes H_{CP}$, the smallest group that can reproduce the NuFIT data is $\Delta(24)\cong S_{4}$ with $n=2$. In case of $(\varphi_{1},\varphi_{2})=\left(-\frac{\pi}{2},\,\frac{\pi}{2}\right)$, the pattern $U^{I}$ reduces to the TM1 form [103, 104] with its first column fixed as $\left(-\frac{2}{\sqrt{6}},\,-\frac{1}{\sqrt{6}},\,\frac{1}{\sqrt{6}}\right)^{T}$, as can seen from Eq. (3.2), and it agrees with the NuFIT limits for both NO and IO. The TM2-type patterns, namely $U^{II}$ with $(\varphi_{1},\varphi_{2})=(0,\pi)$ and $U^{V}$ featuring the trimaximal column $\left(\frac{1}{\sqrt{3}},\,\frac{1}{\sqrt{3}},\,\frac{1}{\sqrt{3}}\right)^{T}$ (see Eq. (3.14) and Eq. (3.36)), don’t agree with JUNO’s latest results at $3.6\sigma$ level, although it is close to the $3\sigma$ upper limit of NuFIT for both NO and IO. Among the one-parameter patterns originating from $A_{5}\rtimes H_{CP}$, the GR2 structure $U^{III}$—which preserves the second column of the Golden Ratio mixing pattern, $\left(\sqrt{\frac{1}{\sqrt{5}\,\phi_{g}}},\,\sqrt{\frac{\phi_{g}}{2\sqrt{5}}},\,\sqrt{\frac{\phi_{g}}{2\sqrt{5}}}\right)^{T}$ (see Eq. (3.25))—remains compatible with the NuFIT constraints in both NO and IO. By contrast, the real mixing pattern $U^{IV}$, characterized by the fixed first column $\left(\frac{\phi_{g}}{2},\,\frac{\phi_{g}-1}{2},\,\frac{1}{2}\right)^{T}$ (see Eq. (3.31)), is only viable in NO.
@@
We display in figure 1 the best-fit values of the three mixing angles and $\delta_{CP}$ for all viable one-parameter models under both NO and IO neutrino mass spectrum. The dashed lines indicate the central values of the observables from the NuFIT global fit. The light-red band in the $\sin^{2}\theta_{12}$ panel corresponds to the $3\sigma$ interval derived from the recent JUNO measurement, see Eq. (5.4), while the darker and narrower red band indicates the projected $3\sigma$ precision after six years of data taking, assuming the current JUNO central value $\sin^{2}\theta_{12}=0.3092$ and the prospective $1\sigma$ uncertainty of $0.5\%$ [34]. For NO, three one-parameter mixing patterns — $U^{I}$ (TM1), $U^{IV}_{1}$ and $U^{IV}_{2}$ — yield predictions for $\sin^{2}\theta_{12}$ that fall within the current JUNO constraint. In contrast, the patterns $U^{II}$ (TM2), $U^{III}$ (GR2), and $U^{V}$ (TM2) are disfavored by the JUNO data at the $3\sigma$ level. None of the one-parameter patterns, however, would remain compatible with the much better sensitivity expected after six years of operation if the current best fit value would not change significantly. For IO, only the pattern $U^{I}$ (TM1) lies in the present JUNO $3\sigma$ allowed interval, and all patterns would be excluded once the planned future precision is reached. We also observe that the predicted values of $\sin^{2}\theta_{13}$ for nearly all patterns lie very close to the NuFIT best-fit line, reflecting the excellent experimental precision already achieved for this observable.
Finally, the red bands shown for $\sin^{2}\theta_{23}$ and $\delta_{CP}$ represent the anticipated future $3\sigma$ sensitivities of next-generation long-baseline experiments such as DUNE [95] and T2HK [96]. As we have discussed, the future experimental uncertainties on $\sin^{2}\theta_{23}$ and $\delta_{CP}$ foreseen to be achieved by T2HK and DUNE depend significantly on the true (i.e., best fit) values of $\sin^{2}\theta_{23}$ and $\delta_{CP}$. For the present NuFIT best-fit values, $\sin^{2}\theta_{23}=0.47$ (NO) and $\sin^{2}\theta_{23}=0.55$ (IO), the prospective $1\sigma$ uncertainty is approximately $3\%$, which we adopt as a representative estimate in the present study, as we have indicated. In the case of $\delta_{CP}$, as explained earlier, we adopt $12^{\circ}$ as prospective $1\sigma$ uncertainty in our analysis.
Combining the numerical results in table 4, we observe that the patterns $U^{I}$, $U^{II}$, and $U^{III}$ predict maximal atmospheric neutrino mixing and CP-violation, $\theta_{23}=\pi/4$ and $\delta_{CP}=3\pi/2$. In contrast, the real mixing patterns $U^{IV}_{1}$ and $U^{IV}_{2}$ imply CP conservation, but with different values of $\delta_{CP}=0$ and $\pi$. Neither maximal CP-violating nor CP-conserving $\delta_{CP}$ is predicted by the $U^{V}_{1}$ and $U^{V}_{2}$ patterns, for which also $\theta_{23}\neq\pi/4$. Hence, the precise measurements of $\theta_{23}$ and $\delta_{CP}$ will play an important role in discriminating between these viable mixing schemes.
@@
For $\sin^{2}\theta_{12}$, the predicted profiles are exceedingly narrow because $\sin^{2}\theta_{12}$ is tightly constrained by the small uncertainties in $\sin^{2}\theta_{13}$, as implied by Eqs. (3.7), (3.28), (3.34) and (3.41) for the $U^{I}$, $U^{III}$, $U^{IV}_{1}$ and $U^{V}_{1}$ patterns, respectively. In the NO case, the prediction of $U^{V}_{1}$ for $\sin^{2}\theta_{12}$ is disfavored by the latest JUNO data, whereas $U^{I}$ and $U^{IV}_{1}$ remain viable at $3\sigma$. In the IO case, $U^{III}$ is excluded by the current JUNO data, while $U^{I}$ is consistent with the data. If JUNO achieves the projected $0.5\%$ precision and the current best fit values would not change significantly, all these patterns could be tested in both NO and IO cases. These conclusions regarding $\sin^{2}\theta_{12}$ for the one-parameter mixing patterns are fully consistent with the earlier analysis based on the best-fit results.
For $\sin^{2}\theta_{23}$, the patterns $U^{I}$ and $U^{III}$ predict $\sin^{2}\theta_{23}=0.5$, while $U^{IV}_{1}$ and $U^{V}_{1}$ yield sharply peaked predictions with $\sin^{2}\theta_{23}>0.5$, due to correlations with $\sin^{2}\theta_{13}$ encoded in Eqs. (3.34) and (3.41). Taking the future projected sensitivities into account, the predicted ranges of $U^{IV}_{1}$ and $U^{V}_{1}$ could be probed in the NO case, while in the IO case both $U^{I}$ and $U^{III}$ could be tested. Regarding the Dirac phase, $U^{I}$ and $U^{III}$ predict $\delta_{CP}=1.5\pi$, while $U^{IV}_{1}$ predicts $\delta_{CP}=\pi$. i.e., absence of CP-violation in neutrino oscillations. In the NO case, the prediction $\delta_{CP}=1.5\pi$ lies outside the projected $3\sigma$ allowed range of the prospective DUNE and T2HK data, whereas $\delta_{CP}=\pi$ remains within it. In contrast, for the IO case, $\delta_{CP}=1.5\pi$ falls within the $3\sigma$ allowed interval.
@@
In summary, the JUNO 59.1-day data imposes significant restrictions on the landscape of one-parameter mixing patterns. The $3\sigma$ interval of $\sin^{2}\theta_{12}$ from JUNO excludes several patterns that would otherwise remain compatible with the NuFIT global fit. We have focused on the flavour symmetry groups $A_{5}$, $\Sigma(168)$ and $D(6n^{2})$ with $n\leq 4$. For NO, we find that only $U^{I}$ (TM1), originating from $S_{4}\rtimes H_{CP}$, and $U^{IV}_{1,2}$, originating from $A_{5}\rtimes H_{CP}$, predict $\sin^{2}\theta_{12}$ values consistent with the current JUNO constraint. For IO, the viable space shrinks further, leaving $U^{I}$ (TM1) as the sole surviving option. Patterns such as $U^{II}$ (TM2), $U^{III}$ (GR2) and $U^{V}$ (TM2), although acceptable under NuFIT alone, become disfavored once the JUNO likelihood is taken into account. It should be emphasized that, for the mixing pattern $U^{I}$, additional phenomenologically viable textures—compatible with the current JUNO data—will emerge other than the TM1 form once higher values of $n$ in the group $\Delta(6n^{2})$ are considered. The future precision measurements of $\sin^{2}\theta_{12}$, $\sin^{2}\theta_{23}$, and $\delta_{CP}$ will provide additional discriminating power for testing these one-parameter mixing patterns.
@@ 5.2 Lepton mixing matrix depending on two free parameters
For the two-parameter mixing patterns derived from $A_{5}\rtimes H_{CP}$, we find 10 and 4 phenomenologically viable realizations of $U^{VI}$ for NO and IO, respectively, as listed in table 5. The patterns $U^{VI}_{2}$, $U^{VI}_{4,1}$, $U^{VI}_{4,2}$, $U^{VI}_{4,3}$, $U^{VI}_{4,4}$, $U^{VI}_{6,1}$ and $U^{VI}_{6,2}$ predict CP-conserving Dirac and Majorana phases, consistent with the general argument presented in section 4.1. Consequently, these patterns are compatible with present data in the NO case, while they are disfavored for IO mass spectrum. By contrast, $U^{VI}_{1}$, $U^{VI}_{3,2}$ and $U^{VI}_{5}$ yield nontrivial CP-violating phases and can accommodate the current experimental data for both NO and IO. $U^{VI}_{3,1}$ can provide viable predictions only for the IO scenario. Under the sum rule constraints in Eq. (4.7), the allowed regions of the lepton observables given in table 10 become narrower than the corresponding $3\sigma$ intervals in table 3. The likelihood profiles for $\sin^{2}\theta_{12}$, $\sin^{2}\theta_{23}$ and $\delta_{CP}$ are shown in figure 4 for selected viable patterns $U^{VI}$ with $\chi^{2}_{\min}\leq 9$. The patterns $U^{VI}_{4,1}$ and $U^{VI}_{6,2}$ for NO, as well as $U^{VI}_{3,1}$, $U^{VI}_{3,2}$ and $U^{VI}_{5}$ for IO, are not shown in the figure.
@@
In the case of the predictions for $\sin^{2}\theta_{12}$, figure 4 shows that the patterns $U^{VI}_{4,2}$, $U^{VI}_{4,4}$, and $U^{VI}_{6,1}$ are consistent with the NuFIT data at the $2\sigma$ level, whereas $U^{VI}_{1}$, $U^{VI}_{2}$, $U^{VI}_{3,2}$, $U^{VI}_{4,3}$, and $U^{VI}_{5}$ remain compatible at $3\sigma$ for the NO scenario. In case of IO, only $U^{VI}_{1}$ is consistent at $2\sigma$. The patterns $U^{VI}_{1}$ and $U^{VI}_{2}$ yield very sharp predictions for $\sin^{2}\theta_{12}$, since $\sin^{2}\theta_{12}$ is correlated with $\sin^{2}\theta_{13}$, see Eq. (4.7). The predictions of the patterns $U^{VI}_{1}$, $U^{VI}_{2}$, $U^{VI}_{3,2}$, and $U^{VI}_{5}$ partially overlap with the $3\sigma$ range of the JUNO 59.1-day measurement, but they are disfavored at the $2\sigma$ level and are expected to be excluded once JUNO reaches its projected six-year precision of $0.5\%$. The remaining four cases $U^{VI}_{4,2}$, $U^{VI}_{4,3}$, $U^{VI}_{4,4}$ and $U^{VI}_{6,1}$ are consistent with both the current and projected JUNO sensitivities for both the NO and IO scenarios.
@@
In what concerns the predictions for $\sin^{2}\theta_{23}$, the patterns $U^{VI}_{3,2}$, $U^{VI}_{4,2}$, $U^{VI}_{4,3}$ and $U^{VI}_{5}$ could be excluded by the projected DUNE and T2HK sensitivities in the NO case, as their predicted values of $\theta_{23}$ lie in second octant. For IO, the pattern $U^{VI}_{1}$ would remain consistent with the data with the anticipated experimental precision. Moreover, all patterns that predict CP-conserving Dirac phase can be tested by future measurements, whereas those yielding nontrivial CP-violating $\delta_{CP}$ would remain viable according to the considered projected sensitivities.
@@
For the two-parameter mixing patterns originating from $\Sigma(168)\rtimes H_{CP}$, we identify 5 viable realizations for NO and 2 for IO, as summarized in table 5. In the NO case, given the fact that $U^{VII}_{2,i}$ is a real matrix (see Eq. (4.14)), the four patterns $U^{VII}_{2,i}$ with $i=1,2,3,4$ predict CP-conserving Dirac phase — 0 by $U^{VII}_{2,1}$ and $U^{VII}_{2,3}$ and and $\pi$ by $U^{VII}_{2,2}$ and $U^{VII}_{2,4}$. The pattern $U^{VII}_{1,1}$ instead yields CP-violating Dirac phase and it is compatible with global oscillation data for both mass orderings, whereas $U^{VII}_{1,2}$ is viable only for IO. Figure 4 displays the likelihood profiles of 3 patterns $U^{VII}_{2,2}$, $U^{VII}_{2,3}$ and $U^{VII}_{2,4}$. All of them predict $\sin^{2}\theta_{12}$ values lying within both the current JUNO 59.1-day and the projected six-year precision $3\sigma$ intervals. Their predictions for $\sin^{2}\theta_{23}$, however, show different levels of compatibility with future accelerator experiments sensitivities provided the current NuFIT best fit value including the SK data would not change drastically: while the $U^{VII}_{2,2}$ and $U^{VII}_{2,4}$ predictions fall outside the expected DUNE and T2HK ranges, the allowed region of $U^{VII}_{2,3}$ partially overlaps with the projected one. Since all these patterns predict CP-conserving Dirac phase, $\delta_{CP}=0$ or $\pi$, they will be critically tested once DUNE and T2HK achieve their anticipated precision on $\delta_{CP}$.